If the sides of a triangle are 3, 4, and 5, then, to the nearest degree, the measure of the smallest angle of the triangle is....
Answer by sudhanshu_kmr(1152)


cos@ = 4^2+5^2- 3^2 / (2*4*5)
= 30/40
= 3/4
@ = 41.409